the temporary tire forced us to decrease our average speed by 15 mph. if the 39 mile drive usually takes 45 minutes how much time should i budget for the drive home
Answer by fcabanski(1391)

D = rt


39=r*3/4


39*4/3 = r = 52 mph


52-15 = 37 mph


D=rt ---> D/r = t


39/37 = 1.054054... = 1 hour 3.24 minutes